Dixboro Company manufactures one product and accounts for costs using a job cost system. You have obtained the following information from the corporation’s books and records for the fiscal year ended May 31, year 1:

  • Total manufacturing cost during the year was $16,500,000 based on actual direct material, actual direct labor, and applied manufacturing overhead.
  • Manufacturing overhead was applied to work in process at 150 percent (1.5 times) direct labor dollars. Applied manufacturing overhead for the year was 30 percent of the total manufacturing cost during the year.

Required:

Compute actual direct material used, actual direct labor, and applied manufacturing overhead. (Hint: The total of these costs is $16,500,000.) (Round intermediate calculations to 2 decimal places.)

Southern Rim Parts estimates its manufacturing overhead to be $490,500 and its direct labor costs to be $1,090,000 for year 1. The first three jobs that Southern Rim worked on had actual direct labor costs of $68,000 for Job 301, $93,000 for Job 302, and $190,000 for Job 303. For the year, actual manufacturing overhead was $384,000 and total direct labor cost was $831,000. Manufacturing overhead is applied to jobs on the basis of direct labor costs using predetermined rates.


Required:

a. How much overhead was assigned to each of the three jobs, 301, 302, and 303?

b. What was the over- or underapplied manufacturing overhead for year 1?

For which of the following businesses would a job costing system be appropriate?

Multiple Choice

  • Auto repair shop.

  • Crude oil refinery.

  • Drug manufacturer.

  • Root beer producer.
Answer #1
Solution-1
Let Manufacturing Overhead Applied= X
Manufacturing Overhead Applied = 150% X Direct Labour
Manufacturing Overhead Applied= 30%X Total Manufacturing Cost
150%X Direct Labour=30% X total Manufacturing cost
Total Manufacturing Cost=$16,500,000
150% DL= 30%*16500000
Direct labor =$3300000
Manufacturing OH Apllied= $3300000X 150%=$4950000
Direct Material Used = Total Manufacturing Cost- Direct Labor- Manufacturing OH Applied
=16500000-3300000-4950000=$8250000
Solution-2
Predetermined OH Rate= Estimated Manufacturing OH/ Direct Labour Cost'
=$490500/1090000=45%
a. Overhead Assigned to Job-301, 302,303
Job 301= 68000X45%=30600
Job 302=93000X45%=41850
Job 303= 190000X45%=85500
b. Computation of Overapplied / Under Applied OH
Actual Overhea dincurred:= 384000
Less: Appled Overhead= 157950
( 30600+41850+85500)
Under applied Overhead=226050